OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, 9).

OpenStudy (campbell_st):

ok... so the focus is above the vertex so the parabola is concave up. do you know the focal length...?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

jello?

OpenStudy (campbell_st):

ok... so a vertex form for the parabola is (x - h)^ = 4a(y - k) (h, k) is the vertex and a is the focal length you know the vertex is (0,0) so you have x^2 = 4ay substitute the focal length to get the equation

OpenStudy (anonymous):

but that wasn't an option

OpenStudy (anonymous):

jk... so the answer would be.....1/36x^2=y? @campbell_st

OpenStudy (campbell_st):

yes, thats my best guess.. well its not really a guess

OpenStudy (anonymous):

thank you!

OpenStudy (campbell_st):

glad to help